OpenStudy (cameron2000):

WILL FAN AND MEDAL PLEASE HELP!!!!!!!!!!!!!!! QUICK!!!!!!! Find the quotient of quantity of 6 times x to the 3rd power minus 18 times x to the 2nd power minus 12 times x all over negative 6 times x. −x2 − 3x − 2 x2 − 18x − 12 −x2 + 18x − 12 −x2 + 3x + 2

OpenStudy (cameron2000):

Will fan and medal please help

OpenStudy (cameron2000):

@.Sam.

OpenStudy (cameron2000):

@jigglypuff314

OpenStudy (cameron2000):

@jigglypuff314

OpenStudy (cameron2000):

@.Sam.

imqwerty (imqwerty):

okay first we write the given information in mathematical terms- we have this-> Find the quotient of quantity of 6 times x to the 3rd power minus 18 times x to the 2nd power minus 12 times x all over negative 6 times x. \(\Large \frac{6 \times x^3-18 \times x^2-12 \times x}{-6 \times x}\) can you simplify it from here? =]

OpenStudy (cameron2000):

dont you switch everything to the other row if it is negative?

imqwerty (imqwerty):

yes true we can also cancel out the common terms from the numerator and the denominator

OpenStudy (cameron2000):

ok so it would be B?

imqwerty (imqwerty):

umm nope we can start like this- \(\Large \frac{6 \times x^3-18 \times x^2-12 \times x}{-6 \times x}\) \(\Large \frac{6x(1x^2-3x-2)}{-6x}\) here we cancel out 6x from numerator and denominator to get this- \(\Large (-x^2+3x+2)\)

OpenStudy (cameron2000):

oh ok, i understand now

imqwerty (imqwerty):

okay =]

OpenStudy (cameron2000):

thank you

imqwerty (imqwerty):

np